how many ounces of 40% acid and 80% acid must be mixed together to get 160ounces of a 70% acid
Answer by oberobic(2304)

The first rule is to keep track of how much pure stuff you need.
160 oz of 70% acid would have 160*.7 = 112 oz. of pure acid.
.
x = oz of 40% acid to mix
.
y = oz of 80% acid to mix
.
x+y = 160
so
y = 160-x
.
.4(x) + .8(160-x) = 112
.
Multiply by 10 to eliminate decimals.
.
4x +8(160-x) = 1120
.
4x +1280 -8x = 1120
.
-4x = -160
.
x = 40
.
y = 160-x
.
y = 120
.
Answer: 40 oz of 40% acid + 120 oz of 70% solution
.
Check to see if this answer produces the right quantity of pure stuff in the solution.
.4*40 + .8*120 = 112 oz
Correct.
.
Done.